# GenRecal: Generation after Recalibration from Large to Small Vision-Language Models

### v2

@@ Abstract
00footnotetext: * Work Done during Internship.
Figure 2: (Left) Comparison of the challenging benchmark performances, MMB [liu2023mmbench], $M$-Vet [yu2023mm], MMMU [yue2023mmmu], and MMMU-Pro [yue2024mmmu] by changing large VLMs. The more powerful large VLMs we select, the greater the performance improvement we can achieve. (Right) Comparing the performance of the challenging benchmark: MMMU [yue2023mmmu], with GenRecal and various VLMs across model sizes. Note that all the experiments in Fig. 1 and Fig. 2 are conducted on the equal training dataset.
@@ 1 Introduction
However, the increasing scale of recent VLMs introduces substantial computational overhead, limiting their practicality in real-world scenarios—particularly for on-device deployment. To mitigate this challenge, recent work has explored distillation techniques [cai2024llava, shu2024llava, feng2024align] that transfer knowledge from large (teacher) VLMs to smaller (student) ones. Yet, existing distillation methods suffer from a fundamental limitation: they typically assume that teacher and student produce sequences of equal token length, enabling token-level distance metrics such as KL divergence. This assumption breaks down when the two models use different tokenizers or adopt different image-splitting mechanisms (see Appendix A), making the distillation process infeasible. Consequently, the range of viable teacher–student VLM pairs becomes severely restricted (see Appendix B). In other words, current distillation approaches only operate when the teacher and student share identical vocabulary sizes, token split, and token index ordering schemes (referred to as token types), as illustrated on the left side of Fig. 1. These differences naturally lead to mismatched input or output token lengths across VLMs, preventing effective distillation. Note that even VLMs within the same family may employ different token types, further hindering compatibility.
Figure 3: Overview of the GenRecal architecture and its training stages. We denote the question and answer tokens from the small VLM as $q_{s}$ and $a_{s}$, and those from the large VLM as $q_{l}$ and $a_{l}$. For simplicity, the vision encoder and image-token projector are omitted from the illustration. Note that Recalibrator is used only during training as a bridge between heterogeneous VLMs. During inference, Recalibrator is removed, leaving the small VLM’s architecture unchanged and adding no extra computational cost.
@@
To demonstrate the necessity of GenRecal, we first compare its performance with that of traditional distillation, as shown in experiment types C and D of Fig. 1. Although both experiments employ the same token types of VLMs—Qwen2-VL-72B as the teacher and Qwen2-VL-7B as the student—GenRecal consistently outperforms the traditional distillation method implemented by LLaVA-KD [cai2024llava]. This result highlights that aligning feature representations between large and small VLMs is crucial for minimizing information loss during distillation, even when the models are homogeneous. Moreover, when we replace the large VLM from Qwen2-VL-72B with the more powerful InternVL2.5-78B, GenRecal achieves enhanced performance. Subsequently, replacing the small VLM with InternVL2.5-8B leads to an even stronger model, as summarized on the right side of Fig. 1. These experiments highlight the significance of GenRecal because traditional methods cannot handle VLM pairs with different token types.
@@ 2 Related Work
Figure 4: Overview of regularization simultaneously done with first training stage-alignment.
3 GenRecal: Generation after Recalibration
3.1 Model Architecture & Various Accommodation
We briefly illustrate the overall model architecture and training procedure in Fig. 3. GenRecal consists of three main components: a large (teacher) VLM, a small (student) VLM, and a Recalibrator module. For the large VLMs, we select models with over 72B parameters to ensure competitive performance: NVLM-72B (Qwen2-72B) [nvlm2024], Qwen2-VL-72B (Qwen2-72B) [wang2024qwen2vl], InternVL2-76B (Llama3-70B) [chen2024far], and InternVL2.5-78B (Qwen2.5-72B) [chen2024expanding]. For the small VLMs, we adopt various model sizes, including Qwen2-VL-2B/7B (Qwen2-2B/7B) and InternVL2.5-1B/2B/4B/8B (InternLM2.5-1.8B/7B and Qwen2.5-0.5B/3B/72B). Note that the parentheses denote the LLMs used within each VLM. Throughout this section, we divide a VLM architecture into four key modules: the vision encoder, vision projector, VLM-body, and VLM-head (i.e. language head). Recalibrator is designed to acquire shared feature representations within a common latent space. It consists of two decoder blocks and two projection layers, as depicted in Fig. 3(c). We denote the decoder blocks as Rec-body, and the projection layers as Rec-proj-pre and Rec-proj-post. Rec-body adopts the same structural design as the student VLM’s decoder, while both Rec-proj-pre and Rec-proj-post are implemented as single linear layers.
3.2 Design and Propagation of Recalibrator
To construct a shared representation space, we first extract and align hidden features from both the large and small VLMs. Since the hidden space of a large VLM typically encompasses a richer vocabulary and more expressive representations, we define it as the shared latent space which the small VLM is projected into. However, direct alignment of the two VLMs’ tokens within this space is infeasible due to discrepancies in vocabulary size, token split, and token index ordering. To overcome this challenge, we draw inspiration from the principle of autoregressive modeling, wherein a model predicts answer tokens for the given question tokens.
$=\textit{VLM-body}_{l}([q_{l},a_{l}]),\quad$ $(l:\text{large})$ (1)
where $q$ and $a$ denote the question and answer tokens produced by each model’s vision encoder, vision projector, and word embedding [mikolov2013efficient]. We then extract the question features $z_{q_{s}}$ from the small VLM and the answer features $z_{a_{l}}$ from the large VLM, and concatenate them to form a joint sequence: $z_{q_{s}},z_{a_{l}}]$, which is then passed through Recalibrator. Because directly embedding both features into a shared space is infeasible due to different token types, we instead design an autoregressive loss that predicts the answer token index of the large VLM given the question token index of the small VLM. This objective encourages Recalibrator to effectively project the small VLM’s features into the large VLM’s latent space.
@@ 3.2 Design and Propagation of Recalibrator
However, the hidden dimensions of large and small VLMs often differ. To address this, in Recalibrator, we first apply Rec-proj-pre to $z_{a_{l}}$, aligning its dimensionality to that of $z_{q_{s}}$. The concatenated sequence $z_{q_{s}},\textit{Rec-proj-pre}(z_{a_{l}})]$ is then propagated through Rec-body decoder blocks. The output of Rec-body is subsequently passed through Rec-proj-post to restore the hidden dimensionality of the large VLM. Finally, the answer part of the resulting features $r_{q_{s}},r_{a_{l}}]=\textit{Recalibrator}([z_{q_{s}},z_{a_{l}}])$ are fed into VLM-head of the large VLM to compute the autoregressive loss:
$\mathcal{L}_{ar}=\text{CE}(\textit{VLM-head}_{l}(r_{a_{l}}),gt_{l}),$ (2)
where $gt_{l}$ means answer token index from large VLM. Additionally, we employ a KL divergence objective (see Algo. 1 in Appendix E) to further enhance Recalibrator’s projection capability. This process enables the large VLM to interpret the hidden features of the small VLM, thus establishing a bridge for general-purpose distillation across heterogeneous token types.
@@
3.3 Training Process
To achieve general-purpose distillation, we adopt a three-stage training process. In the first stage, only Recalibrator is trained while keeping all parameters of both the large and small VLMs frozen, as illustrated in Fig. 3(a). During this stage, we compute both the autoregressive loss and the KL divergence. This step primarily serves to align the feature representations between the large and small VLMs. The detailed procedures for computing the loss functions are presented in Algo. 1 in Appendix E, where the purple components indicate the trainable parameters.
An important finding in this stage is the effectiveness of a regularization term (see Fig. 4), which plays a crucial role in preventing the feature representations from drifting too far from those of the large VLM. Without this regularization, the final distillation performance drops significantly, as shown in the experimental section. This degradation occurs due to insufficient feature alignment between the models. Therefore, we introduce this regularization term to better harmonize the feature representations of the large and small VLMs. The detailed implementation of this regularization is provided in Algo. 2 in Appendix E. Note that the first stage jointly optimizes both objectives from Algo. 1 and 2.
In the second stage (Algo. 3 in Appendix E), we retain the same loss functions from the first stage and additionally incorporate the small VLM’s own autoregressive loss. As depicted in Fig. 3(b), this stage enables knowledge distillation from the large VLM to the small one by training the small VLM’s VLM-body to adapt to the shared feature representation and acquire richer knowledge. In the final stage, we remove both Recalibrator and the large VLM, and fine-tune the student VLM, allowing it to perform inference independently. Specifically, we train all parameters of the student VLM except for the vision encoder through supervised fine-tuning (SFT) to further enhance its instruction-following capability.
4 Experiments
Figure 6: An overview of our training pipeline, illustrating both the question prompt and the measurement/legend annotations (top), followed by t-SNE visualizations (bottom) of teacher and student VLM pairings at the initial and final training stages. The question prompt (upper-left) shows the format of the question, while the measurement and legend box (upper-right) shows key model components to measure. Each scatter plot in the lower panels corresponds to a different combination of teacher and student VLM sizes, capturing how the learned representations evolve from early to later training iterations.
4.1 Implementation Detail
To ensure successful reproducibility, we outline three key technical aspects and details of GenRecal: (a) the structure detail of Recalibrator and (b) specific details of training and evaluation.
(a) Structure Detail of Recalibrator. This is composed of two transformer decoder blocks (Rec-body) and two projectors (Rec-proj-pre and Rec-proj-post). The decoder blocks follow all configurations of small VLMs’ decoder block such as causal mask, hidden dimensions, number of heads, and FFN structure. In addition, the projectors include the single layer linear module. For propagation in Recalibrator, there are some critical techniques that should be considered: new positional embedding (NPE) should be assigned to input sequence for Recalib-body. This is because the input sequence is based on question tokens in small VLMs and answer tokens in large VLMs, so they naturally have different type of positional embeddings and different position-ids as well. Therefore, we employ another RoPE [su2024roformer] to realign their positional embeddings, and their position-ids are re-assigned as well. Lastly, we apply additional layer-norm [ba2016layernormalization] into the output features of Recalibrator for stable adaptation, before meeting VLM-head of large VLM.
@@ 4.1 Implementation Detail
4.2 Validation and Analysis of GenRecal
Higher Size, Higher Performance. Throughout its realization of GenRecal, Tab. 1 demonstrates that GenRecal outperforms not only previous standard model size VLMs but also smaller size VLMs, highlighting that GenRecal generalizes well across different model sizes. From this table, we can say that if we choose more high-performing small VLMs like InternVL2.5-8B [chen2024expanding] significantly impacts the overall distillation results as well. Furthermore, when we choose same large VLMs, selecting more smaller size VLMs makes the lower distillation performances in the end. It implies that employing more capable small VLMs is crucial for achieving higher distillation performances.
@@ 4.2 Validation and Analysis of GenRecal
Figure 7: Depicting the process of (a) computing cosine similarity, (b) the result without regularization or (c) with regularization. Because the number of the embedded tokens for small and large VLMs are naturally different due to vocabulary size, token split, and token ordering, therefore we do average the output tokens and compute cosine similarity.
Next, we examine the effect of changing large VLMs to determine whether the aforementioned property holds consistently across different large VLMs. As shown in Fig. 2, we can see that the results remain consistent, indicating that selecting more powerful large VLMs leads to greater performance improvements. To further investigate it, we utilize InternVL2.5 series [chen2024expanding] and conduct experiments with various combinations of large and small VLM sizes, as illustrated in Fig. 5. The results also consistently indicates that choosing larger size teacher and student VLMs eventually leads to higher distillation performances.
@@
Importance of Regularization. We conduct an ablation study that investigates how important the regularization term is for the final distillation performances. We first remove the regularization term and analyze the differences before and after its removal. To illustrate this in detail, we simply prepare 10 different question-answer pairs and input them through VLM-body of both small and large VLMs. Afterwards, we propagate their output features into Recalibrator, as described in Fig. 7(a). We then compute a $10\times 10$ cosine similarity matrix between Recalibrator’s output features from the small and large VLMs. To ensure reliable cosine similarity values, we use all training dataset samples and compute the average of the cosine similarity values for all the samples.

### v3

@@ Abstract
Keywords:
Cross-Tokenizer Model Distillation Efficient AI
@@ 1 Introduction
Figure 2: (Left) Comparison of the challenging benchmark performances, MMB [liu2023mmbench], $M$-Vet [yu2023mm], MMMU [yue2023mmmu], and MMMU-Pro [yue2024mmmu] by changing large VLMs. The more powerful large VLMs we select, the greater the performance improvement we can achieve. (Right) Comparing the performance of the challenging benchmark: MMMU [yue2023mmmu], with GenRecal and various VLMs across model sizes. Note that all the experiments in Fig.˜1 and Fig.˜2 are conducted on the equal training dataset.
@@
However, the increasing scale of recent VLMs introduces substantial computational overhead, limiting their practicality in real-world scenarios—particularly for on-device deployment. To mitigate this challenge, recent work has explored distillation techniques [cai2024llava, shu2024llava, feng2024align] that transfer knowledge from large (teacher) VLMs to smaller (student) ones. Yet, existing distillation methods suffer from a fundamental limitation: they typically assume that teacher and student produce sequences of equal token length, enabling token-level distance metrics such as KL divergence. This assumption breaks down when the two models use different tokenizers or adopt different image-splitting mechanisms (see Appendix A), making the distillation process infeasible. Consequently, the range of viable teacher–student VLM pairs becomes severely restricted (see Appendix B). In other words, current distillation approaches only operate when the teacher and student share identical vocabulary sizes, token split, and token index ordering schemes (referred to as token types), as illustrated on the left side of Fig.˜1. These differences naturally lead to mismatched input or output token lengths across VLMs, preventing effective distillation. Note that even VLMs within the same family may employ different token types, further hindering compatibility.
Figure 3: Overview of the GenRecal architecture and its training stages. We denote the question and answer tokens from the small VLM as $q_{s}$ and $a_{s}$, and those from the large VLM as $q_{l}$ and $a_{l}$. For simplicity, the vision encoder and image-token projector are omitted from the illustration. Note that Recalibrator is used only during training as a bridge between heterogeneous VLMs. During inference, Recalibrator is removed, leaving the small VLM’s architecture unchanged and no extra computational cost.
@@
To demonstrate the necessity of GenRecal, we first compare its performance with that of traditional distillation, as shown in experiment types C and D of Fig.˜1. Although both experiments employ the same token types of VLMs—Qwen2-VL-72B as the teacher and Qwen2-VL-7B as the student—GenRecal consistently outperforms the traditional distillation method implemented by LLaVA-KD [cai2024llava]. This result highlights that aligning feature representations between large and small VLMs is crucial for minimizing information loss during distillation, even when the models are homogeneous. Moreover, when we replace the large VLM from Qwen2-VL-72B with the more powerful InternVL2.5-78B, GenRecal achieves enhanced performance. Subsequently, replacing the small VLM with InternVL2.5-8B leads to an even stronger model, as summarized on the right side of Fig.˜1. These experiments highlight the significance of GenRecal because traditional methods cannot handle VLM pairs with different token types.
@@ 2 Related Work
Figure 4: Overview of regularization simultaneously done with first training stage-alignment. Note that its propagation rule is different with Fig.˜3(c) where Recalibrator takes small VLM’s tokens for the question and large VLM’s tokens for the answer. In contrast, in this figure, only large VLM’s tokens (question and answer) are fed into Recalibrator. Basically, Rec-body has the hidden dimension of the small VLM, so we should change this hidden dimension of large VLM’s tokens to that of small VLM’s ones via Proj-pre when the large VLM’s tokens are fed into Recalibrator.
@@
3 GenRecal: Generation after Recalibration
3.1 Model Architecture & Various Accommodation
We briefly illustrate the overall model architecture and training procedure in Fig.˜3. GenRecal consists of three main components: a large (teacher) VLM, a small (student) VLM, and a Recalibrator module. For the large VLMs, we select models with over 72B parameters to ensure competitive performance: NVLM-72B (Qwen2-72B) [nvlm2024], Qwen2-VL-72B (Qwen2-72B) [wang2024qwen2vl], InternVL2-76B (Llama3-70B) [chen2024far], and InternVL2.5-78B (Qwen2.5-72B) [chen2024expanding]. For the small VLMs, we adopt various model sizes, including Qwen2-VL-2B/7B (Qwen2-2B/7B) and InternVL2.5-1B/2B/4B/8B (InternLM2.5-1.8B/7B and Qwen2.5-0.5B/3B/72B). Note that the parentheses denote the LLMs used within each VLM. Throughout this section, we divide a VLM architecture into four key modules: the vision encoder, vision projector, VLM-body, and VLM-head (i.e. language head). Recalibrator is designed to acquire shared feature representations within a common latent space. It consists of two decoder blocks and two projection layers, as depicted in Fig.˜3(c). We denote the decoder blocks as Rec-body, and the projection layers as Proj-pre and Proj-post. The decoder blocks, Rec-body, adopt the same structural design as the student VLM’s decoder, while both Proj-pre and Proj-post are implemented as single linear layers.
Algorithm 1 First stage Loss Functions (Alignment)
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), $gt_{l}$ (label index)
[…]
7:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
Algorithm 2 First stage Regularization (Alignment)
1:Input: ($q_{l}$, $a_{l}$), $gt_{l}$ (label index)
[…]
6:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
3.2 Design and Propagation of Recalibrator
To construct a shared representation space, we first extract and align hidden features from both the large and small VLMs. Since the hidden space of a large VLM typically encompasses a richer vocabulary and more expressive representations, we regard it as the shared latent space which the small VLM is projected into. However, direct alignment of the two VLMs’ tokens within this space is infeasible due to discrepancies in vocabulary size, token split, and token index ordering. To overcome this challenge, we draw inspiration from the principle of autoregressive modeling, wherein a model predicts answer tokens for the given question tokens.
$=\textit{VLM-body}_{l}([q_{l},a_{l}]),\quad$ $(l:\text{large})$ (1)
where $q$ and $a$ denote the question and answer tokens produced by each model’s vision encoder, vision projector, and word embedding [mikolov2013efficient]. We then extract the question features $z_{q_{s}}$ from the small VLM and the answer features $z_{a_{l}}$ from the large VLM, and concatenate them to form a joint sequence: $z_{q_{s}},z_{a_{l}}]$, which is then passed through Recalibrator. Because directly embedding both features into a shared space is infeasible due to different token types, we instead design an autoregressive loss that predicts the answer token index of the large VLM given the question token index of the small VLM. This loss encourages Recalibrator to effectively project the small VLM’s features into the large VLM’s latent space.
Algorithm 3 Second stage Loss Functions (Distillation)
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), ($gt_{l}$, $gt_{s}$) (label index)
[…]
8:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
However, the hidden dimensions of large and small VLMs often differ. To address this, in Recalibrator, we first apply Proj-pre to $z_{a_{l}}$, aligning its dimensionality to that of $z_{q_{s}}$. The concatenated sequence $z_{q_{s}},\textit{Proj-pre}(z_{a_{l}})]$ is then propagated through Rec-body decoder blocks. The output of Rec-body is subsequently passed through Proj-post to restore the hidden dimensionality of the large VLM. This forward rule of Recalibrator is illustrated in Fig.˜3(c).
Finally, the answer part of the resulting features $r_{q_{s}},r_{a_{l}}]=\textit{Recalibrator}([z_{q_{s}},z_{a_{l}}])$ are fed into VLM-head of the large VLM to compute the autoregressive loss:
$\mathcal{L}_{ar}=\text{CE}(\textit{VLM-head}_{l}(r_{a_{l}}),gt_{l}),$ (2)
where $gt_{l}$ means answer token index from large VLM. Additionally, we employ a KL divergence loss to further enhance Recalibrator’s projection capability in first training stage and distill the knowledge from the teacher in second training stage. This process enables the large VLM to interpret the hidden features of the small VLM, thus establishing a bridge for general-purpose distillation across heterogeneous token types.
@@ 3.2 Design and Propagation of Recalibrator
3.3 Training Process
To achieve general-purpose distillation, we adopt a three-stage training process. In the first stage, only Recalibrator is trained while keeping all parameters of both the large and small VLMs frozen, as illustrated in Fig.˜3(a). During this stage, we compute both the autoregressive loss and the KL divergence, described in Algorithm˜1 where the purple components indicate the trainable parameters. This step primarily serves to align the feature representations between the large and small VLMs. For the first stage, we introduce regularization loss, as illustrated in Fig.˜4, which is designed to prevent the feature representations of Recalibrator from drifting too far from those of the large VLM. We both use Algorithm˜1 and Algorithm˜2 at the same time for conducting first stage.
In the second stage, we keep using the same loss functions of the first stage and additionally incorporate the small VLM’s own autoregressive loss in Algorithm˜3. As depicted in Fig.˜3(b), this stage enables knowledge distillation from the large VLM to the small one by training the small VLM’s VLM-body. In the final stage, we remove both Recalibrator and the large VLM, and fine-tune the student VLM. Specifically, we train all parameters of the student VLM except for the vision encoder through supervised fine-tuning (SFT) to further enhance its instruction-following capability.
4 Experiments
@@
4.1 Implementation Detail
To ensure successful reproducibility, we outline three key technical aspects and details of GenRecal: (a) the structure detail of Recalibrator and (b) projection layers (c) specific details of training and evaluation.
(a) Structure Detail of Recalibrator. This is composed of two transformer decoder blocks (Rec-body) and two projectors (Proj-pre and Proj-post). The decoder blocks follow all configurations of small VLMs’ decoder block such as causal mask, hidden dimensions, number of heads, and FFN structure. The reason why transformer decoder blocks are used for Rec-body is because we should handle the concatenated student and teacher features in Fig.˜3, and they are sequence that transformer decoder normally treat. Even, these features should preserve token ordering, so we additionally introduce new positional embeddings (NPE) to adapt it. This design enables effective sequential modeling and representation alignment between heterogeneous VLMs. To realize it, we employ additional RoPE [su2024roformer] to realign their positional embeddings, and their position IDs are re-assigned as well. Lastly, we apply additional layer-norm [ba2016layernormalization] into the output features of Recalibrator for stable adaptation.
(b) Projection Layer Design. Two projection layers (Proj-pre and Proj-post) are linear layers introduced for hidden dimension matching. Similar to the role of vision projector that matches the hidden dimensions between a CLIP-like vision encoder and an LLM, Proj-pre matches the teacher features to the student hidden dimension, while Proj-post restores the dimension back to the teacher hidden dimension so that the large VLM-head can be utilized for distillation. This explanation is technically depicted in Fig.˜3(c) and Fig.˜4.
@@ 4.1 Implementation Detail
4.2 Validation and Analysis for GenRecal
Figure 6: An overview of our training pipeline, illustrating both the question prompt and the measurement/legend annotations (top), followed by t-SNE visualizations (bottom) of teacher and student VLM pairings at the initial and final training stages. The question prompt (upper-left) shows the format of the question, while the measurement and legend box (upper-right) shows key model components to measure. Each scatter plot in the lower panels corresponds to a different combination of teacher and student VLM sizes, capturing how the learned representations evolve from early to later training.
Higher Size, Higher Performance. Throughout its realization of GenRecal, Tab.˜1 demonstrates that GenRecal outperforms not only previous standard model size VLMs but also smaller size VLMs, highlighting that GenRecal generalizes well across different model sizes. From this table, we can say that choosing more high-performing small VLMs like InternVL2.5-8B [chen2024expanding] significantly impacts the overall distillation results as well. Furthermore, when we choose same large VLMs, selecting more smaller size VLMs makes the lower distillation performances in the end. It implies that employing more capable small VLMs is crucial for achieving higher distillation performances.
Next, we examine the effect of changing large VLMs to determine whether the aforementioned property holds consistently across different large VLMs. As shown in Fig.˜2 and Tab.˜2, we can see that selecting more powerful large VLMs leads to greater performance improvements. To further investigate it, we utilize InternVL2.5 series [chen2024expanding] and conduct experiments with various combinations of large and small VLM sizes, as illustrated in Fig.˜5. The results also consistently indicates that choosing larger size teacher and student VLMs eventually leads to higher distillation performances.
[…]
@@ 4.2 Validation and Analysis for GenRecal
Figure 7: The process of (a) computing cosine similarity, (b) the result without regularization or (c) with regularization. Because the number of the embedded tokens for small and large VLMs are naturally different due to vocabulary size, token split, and token ordering, therefore we do average the output tokens and compute cosine similarity.
Importance of Regularization. We conduct an ablation study that investigates how important the regularization term is for the final distillation performances. We first remove the regularization term and analyze the differences before and after its removal. To illustrate this in detail, we simply prepare 10 different question-answer pairs and input them through VLM-body of both small and large VLMs. Afterwards, we propagate their output features into Recalibrator, as described in Fig.˜7(a). We then compute a $10\times 10$ cosine similarity matrix between Recalibrator’s output features from the small and large VLMs. To ensure reliable cosine similarity values, we use all training dataset samples and compute the average of the cosine similarity values for all the samples.
With the regularization term included, as shown in Fig.˜7(c), we observe that the diagonal values in the cosine similarity matrix are significantly higher than the off-diagonal values. This implies that the features of small and large VLMs are explicitly shared through Recalibrator, aligned with the results in Fig.˜6 as well. In contrast, when the regularization term is removed, as depicted in Fig.˜7(b), the diagonal values are comparable with the off-diagonal values. This indicates that Recalibrator fails to explicitly align large and small VLMs’ features, preventing feature matching and sharing. As a result, this leads to a significant drop in final distillation performances, as reported in Tab.˜4. Insufficient alignment of large and student VLMs reduces efficient knowledge transfer during distillation. These findings highlight the crucial role of the regularization term in feature alignment and knowledge transfer for general-purpose distillation.
Autoregressive Loss for Representation Alignment. To align the student representations with the teacher’s latent space, GenRecal uses two objectives: KL divergence loss ($\mathcal{L}_{kl}$) and autoregressive loss ($\mathcal{L}_{ar}$). The KL divergence loss performs standard distillation by matching the logit-level distribution of teacher and student. However, matching distributions alone does not guarantee precise alignment at the token level. The autoregressive loss provides explicit hard-target supervision, allowing it to predict the exact ground-truth token sequence. In short, $\mathcal{L}_{kl}$ transfers knowledge at the distribution level, whereas $\mathcal{L}_{ar}$ ensures exact alignment in the token space-level. As shown in Tab.˜5, removing $\mathcal{L}_{ar}$ consistently degrades performance across benchmarks, confirming that autoregressive supervision is essential for effective representation alignment.
Table 5: Importance of the autoregressive loss ($\mathcal{L}_{ar}$) for representation alignment. We compare the baseline VLMs, GenRecal trained without $\mathcal{L}_{ar}$ (✗, using only $\mathcal{L}_{kl}$), and GenRecal trained with $\mathcal{L}_{ar}$ (✓), using InternVL2.5-78B [chen2024expanding] as the teacher VLM.
@@
Qwen2-VL-7B [wang2024qwen2vl] - 77.5 83.0 58.2 83.0 80.5 62.0 54.1 30.5 53.8 68.6 68.5 65.4
InternVL2.5-8B [chen2024expanding] - 84.8 84.8 64.4 84.6 82.6 62.8 56.0 34.3 54.8 69.7 70.1 68.1
Qwen2-VL-7B-GenRecal (InternVL2.5-78B) 66.2 67.5 47.8 63.4 55.9 49.6 45.8 30.5 52.1 65.3 60.9 55.0
InternVL2.5-8B-GenRecal (InternVL2.5-78B) 68.4 70.1 55.2 68.9 62.5 55.4 50.6 35.1 58.7 68.2 64.4 59.8
Qwen2-VL-7B-GenRecal (InternVL2.5-78B) 93.9 95.3 68.8 88.4 87.4 70.4 65.6 49.6 64.3 70.7 79.2 75.8
InternVL2.5-8B-GenRecal (InternVL2.5-78B) 93.0 93.6 74.9 89.5 88.2 73.2 68.1 48.8 65.3 72.3 81.4 77.1
4.3 Ablation Studies: Training Computation and Configurations
In Tab.˜6(a), we compare the model sizes and training FLOPs of the small VLM, large VLM, and Recalibrator to assess whether Recalibrator introduces additional computational overhead during training. The results show that Recalibrator’s FLOPs are substantially lower than those of both the large and small VLMs. Note that Recalibrator is removed at inference time, thus incurring no additional computational cost during inference. We further conduct a series of ablation studies to examine various configurations affecting the distillation performance of GenRecal. As shown in Tab.˜6(b), we vary the decoder depth of Recalibrator from 1 to 20 and select two depths for layer number of Rec-body based on a trade-off between computational efficiency and performance. Additionally, Tab.˜6(c) demonstrates the importance of employing a new positional embedding (NPE) to handle newly introduced position IDs and to integrate RoPE [su2024roformer] from both large and small VLMs, which have different token types.
To analyze the effect of dataset scale, we control the dataset size from 1M to 9M samples, as shown in Tab.˜6(d). The performance improvements between 5M and 9M samples are marginal, suggesting that a 5M dataset is a practical choice for users with limited training resources. Furthermore, Tab.˜6(e) shows that fine-tuning the large VLM on our dataset yields further gains, indicating that the small VLM trained under the same setting has the potential to match or even surpass the large VLM’s performance.
We also compare GenRecal with recent cross-tokenizer distillation methods in Tab.˜6(f). These baselines perform feature-level distillation via cross-token matching in the logit space by using classical method of Wasserstein distance [boizard2025towards] or optimal transport [cui2025multi], whereas GenRecal explicitly introduces Recalibrator to align hidden features for semantic correspondence. As observed, GenRecal achieves a substantial improvement over prior approaches. It will be discussed in the next section. Lastly, Tab.˜6(g)–(i) demonstrate GenRecal’s strong compatibility across different LLMs, VLM Family, and recently released VLMs. In all cases, GenRecal consistently delivers superior performance, validating its general applicability to diverse model architectures.